# Position Coupling: Leveraging Task Structure for Improved Length Generalization of Transformers

### v1

@@ 5 Theoretical Analyses on 1-layer Transformers
For the sake of simplicity of presentation, we consider a Transformer without any normalization layers, as conventionally done in theoretical constructions by previous works (Yun et al., 2020a, b; Awasthi and Gupta, 2023). For the sake of completeness, readers can find a mathematical formulation of the decoder-only Transformer architecture in Appendix C.
@@
We provide a detailed construction in Appendix D. We highlight that our proof is constructive and does not rely on any universal approximation result of neural networks.
Theorem 5.1 shows that a 1-layer 2-head Transformer is sufficient for implementing addition between two exponentially long integers. We remark that this result can be naturally extended to larger architectures with more layers/heads, with the help of residual connections.
@@ 5.1.1 Probing the Attention Patterns in Trained Transformers with Position Coupling
We discover a striking similarity between the attention patterns in our theoretical construction (Theorem 5.1) and those extracted from a Transformer trained with position coupling and a standard optimizer. In particular, the manually constructed attention patterns described in Tables 11 and 17 in Appendix D closely resemble the actual attention patterns in Figure 7.333Note that they match up to matrix transpose, which is due to the difference in the formulations. We were surprised because the similarity was discovered after we came up with the proof of the theorem. Drawn from this discovery, we claim that a Transformer trained with position coupling spontaneously learns two separate components of the addition task: (1) adding two numbers without carries, and (2) predicting the carries.
Let us revisit the example in Figure 3 and consider predicting ‘7’ (position ID 6) as the next token of ‘0’ (position ID 7). Note that the token ‘7’ is the result of combining the digit-wise sum 6+0=6 and a propagated carry 1. To find out the sum without carry, it is enough for the model to attend to the two previous positions with ID 6: tokens ‘6’ and ‘0’. On the other hand, to predict the carry, the model may attend to the three positions with ID 7: tokens ‘5’, ‘4’, and ‘0’. The reason way we should care about ‘0’ is that considering the sum 5+4 (=9) of the two digits in the operands is not sufficient to determine the existence of the carry. By looking at the token ‘0’ in the response (with position ID 7), we can detect that the actual sum in this position is 10 (=5+4+1, where 1 is another carry propagated from the previous position) and hence we need to propagate a carry 1 to the next position (with ID 6).
Now we inspect the aforementioned claim by examining the attention matrices of an actual trained Transformer. In the model, we discover two different patterns of attention matrices, [^4] Note that the attention matrices depicted in Figure 7 are square, lower-triangular (due to causal attention pattern), and row-stochastic (all entries are nonnegative and the sum of each row equals 1). playing distinct roles. The first attention pattern (top of the figure) seems to correspond to the addition without carries: each token in the response (including ‘=’) attends to two positions needed to find out the sum without carry. Conversely, the second attention pattern (bottom of the figure) seems to correspond to the carry prediction: again, each token in the response attends to three positions required to find out the carry.
Remark.   Similarly to our analysis, Quirke et al. (2023) study the attention patterns of a 1-layer 3-head decoder-only Transformer model trained solely on 5-digit addition. They also observe that each head handles different subtasks of addition, such as digit-wise summation and carry detection.
@@
The proof is deferred to Appendix E. According to the proposition above, the 1-layer Transformer without positional encoding will always output the same values starting from the ‘=’ token, provided that the combination of query tokens is identical, even if their order varies. However, the addition task is permutation-sensitive, meaning that the permuted queries may result in different responses. Therefore, the 1-layer Transformer cannot completely solve the task without positional encoding. It is important to note that this result remains unchanged regardless of the input format: neither reversed format nor index hinting provides any benefit. We also highlight that this impossibility result can be extended to any other permutation-sensitive tasks, such as arithmetic tasks and copy/reverse tasks.
Based on this fact, we write code to directly calculate the maximum EM accuracy on the $m$-digit addition task that a 1-layer decoder-only Transformer can achieve (see Appendix E for the code). The accuracies rapidly decrease to zero: $6.2$% for 3-digit addition, $1$% for 4-digit integers, and $0.13$% for 5-digit integers. We leave it for future work to investigate the necessary conditions of the architecture for implementing addition when other positional encoding schemes are employed.
6 Extension to Other Arithmetic Tasks
The possible utilization of position coupling is not limited to adding two decimal integers. In the following subsections, we explore the effectiveness of position coupling for the addition task with more than two operands (Section 6.1) and the multiplication task with the length of the second operand being fixed (Section 6.2).
6.1 Addition Task with Multiple Summands
The position coupling scheme for the vanilla addition task (with two operands) can naturally extend to the addition task with multiple summands: assign position IDs in ascending order from most significant digits to least significant digits for every operand and the response. Here, we focus on the addition of three summands. To the best of our knowledge, there is no prior work considering multiple summands.
Figure 8: Exact-match accuracy (median over 4 runs) for triple addition task, trained on sequences of length 1–40 with position coupling, NoPE, and random-start APE. For further experiment details, refer to Table 6.
Experiments.
@@ 6.2 Position Coupling for $N\times 2$ Multiplication Tasks
In this subsection, we study length generalization on the $N$-digit $\times$ 2-digit multiplication task in terms of the length $N$ of the first operand, while fixing the length of the second operand by 2. Similar tasks have been studied before (Duan and Shi, 2023; Jelassi et al., 2023); Jelassi et al. (2023) investigates $N\times 3$ using an encoder-only model and Duan and Shi (2023) studies $N\times 1$ with an encoder-decoder Transformer architecture. Besides the architectural difference, Jelassi et al. (2023) fail to observe length generalization with RPE and only achieve it by supplementing a small number of long samples to the training set. Furthermore, although Duan and Shi (2023) provides perfect length generalization results even for test samples 10×10$\times$10 × longer than those observed during training, their approach requires a retraining step with hand-crafted bias correction on attention score matrices.
Data Format & Position Coupling.
We reverse and zero-pad the response, setting the length of it as $N+2$. We couple the position starting from the least significant digits of both operands and response, decrementing the ID as we move to their most significant digits: see Figure 9.
Experiments.
The experimental results showcased in Figure 10 verify the efficacy of position coupling compared to NoPE and random-start APE. We observe that a 1-layer model fails even with position coupling, even for training. However, as the depth increases to 2 or more, it immediately becomes capable of length generalization.
Figure 10: Exact-match accuracies for $N\times 2$ multiplication task, trained on sequences of length 1–40.
@@
We defer the proof to Appendix F. This result suggests that the proposed position coupling scheme for the $N\times 2$ multiplication task sufficiently captures the inherent structure of the task, and thus provides the potential for the trained model to generalize across unseen lengths. Also, we believe that Theorem 6.1 is optimal in terms of the number of attention layers, as the depth-1 model exhibits total failure even for in-distribution samples in our experiment.
7 Applying Position Coupling Beyond Arithmetic Tasks
In this section, we investigate the versatility of position coupling beyond arithmetic tasks. We begin by examining copy/reverse tasks in Section 7.1. Then, we investigate the efficacy of position coupling in managing multi-dimensional tasks in Section 7.2.
7.1 Position Coupling for Copy/Reverse Tasks
We consider the copy task and the reverse task, allowing duplicate tokens in the query. In particular, only 10 characters are considered to appear in the query, while the length of the query can be much longer. Notably, the copy task with duplicates is identified by Zhou et al. (2024a) as a difficult task for decoder-only Transformers to length-generalize, whereas the Transformers can easily length-generalize if there are no duplicate tokens.
Data Format & Position Coupling.
We couple the positions in the query and the response by their correspondence. Note that the position ID assigned to the equal token is different for the two tasks because as per our design principle (Section 3), the equal token is grouped to the response tokens and position IDs have to be consecutive numbers within each group.
Figure 11: Illustration of position coupling for copy/reverse tasks.
Experiments.
We compare the performance of position coupling with NoPE and random-start APE. We train the model on lengths 1–40 and evaluate its performance on lengths from 5 to 300, at intervals of 5. While a 1-layer 4-head model is used for the position coupling, we observe that the same architecture fails to memorize training samples for both NoPE and random-start APE. Therefore, we use a 6-layer 8-head model for the latter cases as it is commonly used in the literature (Zhou et al., 2024a).
The experimental results are described in Figures 12 and 13. For both copy and reverse, position coupling exhibits near-perfect accuracy across the entire test length (7.5$\times$ for the trained length). In contrast, NoPE and random-start APE immediately fail to length-generalize.
Figure 12: Exact-match accuracy (median over 4 runs) for copying task, trained on sequences of length 1–40 with position coupling, NoPE, and random-start APE. For further experiment details, refer to Table 7.
Figure 13: Exact-match accuracy (median over 4 runs) for reversing task, trained on sequences of length 1–40 with position coupling, NoPE, and random-start APE. For further experiment details, refer to Table 7.
7.2 Two-dimensional Position Couplings for Minesweeper Generator Task
The tasks considered thus far are 1D, indicating that both the query and response are structured as 1D objects. This leads us to a natural question: Can position coupling still be effective for multi-dimensional tasks? To examine this question, we explore extending position coupling to accommodate a 2D task, where both the query and response are originally 2D objects. Specifically, we define and investigate a new task called the minesweeper generator. Given a rectangular board where each cell is filled with either ‘M’ (mine) or $\ast$’ (empty cell), the task is to generate a new board of the same size, with each cell containing:
@@ 7.2 Two-dimensional Position Couplings for Minesweeper Generator Task
Experiments.   To assess the efficacy of position coupling, we contrast its performance with NoPE. The training samples are designed with the width and height of the board between 5 and 9 inclusively. We allow the width and height to be different for training samples. We evaluate the test performance on a square board with a width between 5 and 14 inclusively. We also employ a 4-layer 8-head model for position coupling and a 6-layer 8-head model for NoPE. In particular, for position coupling, we use the same embedding layer for both position coupling modules, as this approach empirically performs better than using distinct embedding layers for each module (see Section A.2).
The experimental results are described in Figure 15. Position coupling maintains over 98% accuracy until a width of 12 and near 90% accuracy even at a width of 14. In contrast, NoPE fails even for in-distribution samples. One might worry that the generalizable length of 12 seems only slightly higher than the trained length of 9. However, we stress that our query is a 2D board, therefore the actual length generalization is from 81 to 144.
@@
Figure 14: Position coupling for the two-dimensional ‘minesweeper generator’ task. (Left) The idea of assigning coupled position IDs. (Right) The model receives a flattened sequence of input tokens and 2D position IDs.
@@
Figure 15: EM accuracies for minesweeper generator task, trained on sequences of length (5–9)$\times$(5–9).
8 Related Works
Length Generalization on Arithmetic/Algorithmic Tasks.
@@ 8 Related Works
Despite of simplicity of the tasks, arithmetic and algorithmic tasks have received a lot of attention as subjects to study the length generalization of Transformers. Shen et al. (2023) propose “Random Spacing” and “Recursive Scratchpad”, achieving near-perfect generalization from 10-digits to 12-digits addition. Awasthi and Gupta (2023) devise an auxiliary task for the sorting task and achieves significant length generalization through multitask learning. Ruoss et al. (2023) propose Randomized Positional Encodings to address the problem of the appearance of unseen position IDs in test time. Most recently, Zhou et al. (2024b) demonstrate a possibility of extrapolation to the length 100 with training length 1–40 in the addition task by combining appropriate input format and advanced PE, yet they also observe that the performances are not robust and highly depend on the random seeds.
Analyzing Length Generalization in Theoretical Perspectives
@@
9 Conclusion
Limitations & Future Directions.
We intentionally limited ourselves to the tasks with an explicit structure between the tokens in each sequence. This is because we are proposing a method to instill the known structure of the task into a Transformer by training on short sequences. Designing the coupling of positions for tasks whose structure is implicit or black-box (e.g., for general NLP tasks) remains a fascinating next step: we leave the methodology for uncovering hidden structures and autonomously creating appropriate couplings (without manually designing them) for future work.
We also leave two challenging arithmetic tasks to length-generalize for future work. The first is addition with a varying number of summands, i.e., determining if the model can generalize to adding more integers than the trained samples. The second task is multiplication, where the lengths of both operands can vary.
Acknowledgements
HC, JC, and CY acknowledge support from a Google Gift on the research related to Long Context Transformers. The experiments contained in this work were supported in part through the Google Cloud Platform Credit Award.
Appendix A More Experimental Results on Position Couping
@@
A.1 Decimal Integer Addition Task: Maximum Exact-Match Accuracies
For a better comparison with Zhou et al. (2024b), we report the maximum exact-match (EM) accuracies. See Figures 16 and 17.
Figure 16: Ablation on the trained lengths.
Figure 17: Ablation on the number of layers.
A.2 Two-dimensional Task: Sharing the Position Embedding Modules or Not

### v2

@@ 5 Theoretical Analyses on 1-layer Transformers
For the sake of simplicity of presentation, we consider a Transformer without any normalization layers, as conventionally done in theoretical constructions by previous works (Yun et al., 2020a, b; Awasthi and Gupta, 2023). For the sake of completeness, readers can find a mathematical formulation of the decoder-only Transformer architecture in Appendix D.
@@
We provide our proof in Appendix E. We highlight that our proof is constructive and does not rely on any universal approximation result of neural networks.
Theorem 5.1 shows that a 1-layer 2-head Transformer is sufficient for implementing addition between two exponentially long integers. We emphasize that this result can be naturally extended to larger architectures with more layers/heads, with the help of residual connections.
@@ 5.1.1 Probing the Attention Patterns in Trained Transformers with Position Coupling
We discover a striking similarity between the attention patterns in our theoretical construction (Theorem 5.1) and those extracted from a Transformer trained with position coupling and a standard optimizer. In particular, the manually constructed attention patterns described in Tables 11 and 17 in Appendix E closely resemble the actual attention patterns in Figure 7.222Note that they match up to matrix transpose, which is due to the difference in the formulations. Drawn from this discovery, we claim that a Transformer trained with position coupling spontaneously learns two separate components of the addition task: (1) adding two numbers without carries, and (2) predicting the carries.
Let us revisit the example in Figure 2 and consider predicting ‘7’ (position ID 6) as the next token of ‘0’ (position ID 7). Note that the token ‘7’ is the result of combining the digit-wise sum 6+0=6 and a propagated carry 1. To find out the sum without carry, it is enough for the model to attend to the two previous positions with ID 6: tokens ‘6’ and ‘0’. On the other hand, to predict the carry, the model may attend to the three positions with ID 7: tokens ‘5’, ‘4’, and ‘0’. The reason why we should care about ‘0’ is that considering the sum 5+4 (=9) of the two digits in the operands is not sufficient to determine the existence of the carry. By looking at the token ‘0’ in the response (with position ID 7), we can detect that the actual sum in this position is 10 (=5+4+1, where 1 is another carry propagated from the previous position) and hence we need to propagate a carry 1 to the next position (with ID 6).
Now we inspect the aforementioned claim by examining the attention matrices of an actual trained Transformer. In the model, we discover two different patterns of attention matrices, [^3] The attention matrices depicted in Figure 7 are square, lower-triangular (due to causal attention pattern), and row-stochastic (all entries are nonnegative and the sum of each row equals 1). playing distinct roles. The first attention pattern (top of the figure) seems to correspond to the addition without carries: each token in the response (including ‘=’) attends to two positions needed to find out the sum without carry. Conversely, the second attention pattern (bottom of the figure) seems to correspond to the carry prediction: again, each token in the response attends to three positions required to find out the carry.
Remark.   Similarly to our analysis, Quirke and Barez (2024) study the attention patterns of a 1-layer 3-head decoder-only Transformer model trained solely on 5-digit addition. They also observe that each head handles different subtasks of addition, such as digit-wise summation and carry detection.
@@
The proof is deferred to Appendix F. According to the proposition above, the 1-layer Transformer without positional encoding will always output the same values starting from the ‘=’ token, provided that the combination of query tokens is identical, even if their order varies. However, the addition task is permutation-sensitive, meaning that the permuted queries may result in different responses. Therefore, the 1-layer Transformer cannot completely solve the task without positional encoding. It is important to note that this result remains unchanged regardless of the input format: neither reversed format nor index hinting provides any benefit. We also highlight that this impossibility result can be extended to any other permutation-sensitive tasks, such as arithmetic tasks and copy/reverse tasks.
Based on this, we write code to directly calculate the maximum EM accuracy on the $m$-digit addition task that a 1-layer decoder-only Transformer can achieve (see Appendix F for the code). The accuracies rapidly decrease to zero: $6.2$% for 3-digit addition, $1$% for 4-digit integers, and $0.13$% for 5-digit integers. We leave it for future work to investigate the necessary conditions of the architecture for implementing addition when other positional encoding schemes are employed.
6 Applying Position Coupling Beyond Addition Task
To demonstrate the versatility of position coupling, we consider two other tasks in this section: $N\times 2$ multiplication and a two-dimensional (2D) task. Other example tasks (e.g., addition with multiple summands, copy/reverse allowing duplicates) can be found in Appendix B.
@@ 6.1 Position Coupling for $N\times 2$ Multiplication Tasks
Here, we study length generalization on the $N$-digit $\times$ 2-digit multiplication task in terms of the length $N$ of the first operand, while fixing the length of the second operand by 2. Similar tasks have been studied before (Duan and Shi, 2023; Jelassi et al., 2023); we discuss further in Appendix A.
We reverse and zero-pad the response, setting the length of it as $N+2$. We couple the position starting from the least significant digits of both operands and response, decrementing the ID as we move to their most significant digits: see Figure 17 in Section B.5. The experimental results showcased in Figure 8 verify the efficacy of position coupling compared to NoPE and random-start APE. We observe that a 1-layer model fails even with position coupling, even for training. However, as the depth increases to 2 or more, it immediately becomes capable of length generalization.
Figure 8: $N\times 2$ multiplication task, trained on sequences of length 1–40.
@@
We defer the proof to Appendix G. This result suggests that the proposed position coupling scheme for the $N\times 2$ multiplication task sufficiently captures the inherent structure of the task, and thus provides the potential for the trained model to generalize across unseen lengths. Also, we believe that Theorem 6.1 is optimal in terms of the number of attention layers, as the depth-1 model exhibits total failure even for in-distribution samples in our experiment.
6.2 Two-dimensional Position Coupling for Minesweeper Generator Task
Now, we investigate the extension of position coupling for handling a 2D task, where the query and the response are originally 2D objects. In particular, we define and investigate a task we call minesweeper generator. Given a rectangular board where each cell is filled with either ‘M’ (mine) or $\ast$’ (an empty cell), the task is to generate a new board of the same size, having each cell filled with:
@@ 6.2 Two-dimensional Position Coupling for Minesweeper Generator Task
Figure 9: Position coupling for the two-dimensional ‘minesweeper generator’ task. (Left) The idea of assigning coupled position IDs. (Right) The model receives a flattened sequence of input tokens and two-dimensional position IDs.
Experiments.   To assess the efficacy of position coupling, we contrast its performance with NoPE. The training samples are designed with the width and height of the board between 5 and 9 inclusively. We allow the width and height to be different for training samples. We evaluate the test performance on a square board with a width between 5 and 14 inclusively. We also employ a 4-layer 8-head model for position coupling and a 6-layer 8-head model for NoPE. In particular, for position coupling, we use the same embedding layer for both position coupling modules, as this approach empirically performs better than using distinct embedding layers for each module (see Section B.8).
The experimental results are described in Figure 10. Position coupling maintains over 98% accuracy until a width of 12 and near 90% accuracy even at a width of 14. In contrast, NoPE fails even for in-distribution samples. One might be concerned that the generalizable length of 12 seems only slightly higher than the trained length of 9. However, we stress that our query is a 2D board, therefore the actual length generalization is from 81 to 144.
@@
Figure 10: Minesweeper generator task, trained on sequences of length (5–9)$\times$(5–9).
7 Conclusion
Limitations & Future Directions.   We intentionally limited ourselves to the tasks with an explicit structure between the tokens in each sequence. This is because we are proposing a method to instill the known structure of the task into a Transformer by training on short sequences. Designing the coupling of positions for tasks whose structure is implicit or black-box (e.g., for general NLP tasks) remains a fascinating next step: we leave the methodology for uncovering hidden structures and autonomously creating appropriate couplings (without manually designing them) for future work.
We also leave two challenging arithmetic tasks to length-generalize for future work. One is the addition with a varying number of summands, i.e., determining if the model can generalize to summing multiple integers when trained on samples with fewer summands. The second task is multiplication, where the lengths of both operands can vary. Note that our method is further extended to solve these two challenging length generalization problems in a recent work (Cho et al., 2024).
Acknowledgments and Disclosure of Funding
This work was partly supported by a National Research Foundation of Korea (NRF) grant (No. RS-2024-00421203) funded by the Korean government (MSIT), and an Institute for Information & communications Technology Planning & Evaluation (IITP) grant (No.RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST)) funded by the Korean government (MSIT). HC, JC, and CY acknowledge support from a Google Gift on the research related to Long Context Transformers. The experiments contained in this work were supported in part through a Google Cloud Platform Credit Award.
Appendix A Omitted Backgrounds
A.1 Next-token Prediction with Decoder-only Transformers
Figure 11: Schematic of solving an integer addition task instance using next-token prediction with a decoder-only Transformers. BOS/EOS mean beginning-/end-of-sequence tokens, respectively. PAD means a padding token, used for matching the sequence lengths in a single minibatch of sequences. Here we assume a basic input format (plain, no zero-padding), which is different from that we used in our experiment.
A.2 Related Works
Length Generalization in the Addition Tasks.
@@ A.2 Related Works
Lee et al. (2024) observe that reversing the output in the addition task enables the model to learn a simple function. Shen et al. (2023) propose “Random Spacing” and “Recursive Scratchpad”, achieving near-perfect generalization from 10-digits to 12-digits addition. Zhou et al. (2024a) introduce “index hints”, position markers placed in front of each token, in both the input and output of addition tasks. Most recently, Zhou et al. (2024b) demonstrate a possibility of extrapolation to the length 100 with training length 1–40 in the addition task by combining appropriate input format and advanced PE, yet they also observe that the performances are not robust and highly depend on the random seeds.
Length Generalization in the $N\times M$ Multiplication Task ($M$ is fixed).
Jelassi et al. (2023) investigate $N\times 3$ using an encoder-only model and Duan and Shi (2023) study $N\times 1$ with an encoder-decoder Transformer architecture. Besides the architectural difference, Jelassi et al. (2023) fail to observe length generalization with RPE and only achieve it by supplementing a small number of long samples to the training set. Furthermore, although Duan and Shi (2023) provide perfect length generalization results even for test samples 10×10$\times$10 × longer than those observed during training, their approach requires a retraining step with hand-crafted bias correction on attention score matrices.
Analyzing Length Generalization in Theoretical Perspectives.
@@
Appendix B More Applications & Experiments of Position Couping
B.1 Decimal Integer Addition Task: Scale-up to Length of 500
Here, we demonstrate the scalability of our proposed position coupling approach for large lengths of up to 500. Specifically, we again train a depth-1 decoder-only model for the addition task and evaluate the performance for instances with up to 500 digits. The results are shown in Figure 12. We notice that at a train length of 160, we achieve excellent length generalization for 500-digit addition. On the other hand, training on sequences of length up to 40 or 80 is insufficient for extreme length generalization. The results demonstrate that position coupling, as an approach, is highly scalable.
Figure 12: The exact-match accuracies obtained by training a depth-1 transformer on the addition task. We see that while training with sequences of length up to 40 and 80 is insufficient for generalization to large lengths, at training length 160 we achieve strong performance for lengths up to 500. The experimental details can be found in Table 3.
B.2 Decimal Integer Addition Task: Operands of Different Lengths
In the main text, we mainly focus on the test examples of additions where the lengths of both operands are the same. For the sake of completeness, we also report the evaluation results for the cases where the operand lengths can be different (although the zero-padding is applied to ensure the consistency of the data format). See Figure 13.
Figure 13: Exact-match accuracies (%) on additions with operands of different lengths. Different heatmap corresponds to different trained length of operands (1–10, 1–20, 1–30, and 1–40, expressed with a red box for each). For each heatmap, the $x$-axis and the $y$-axis are for the length of the first and the second operand, respectively.
@@ B.2 Decimal Integer Addition Task: Operands of Different Lengths
B.3 Decimal Integer Addition Task: Maximum Exact-Match Accuracies
A prior work by Zhou et al. (2024b) provides a similar analysis on the addition tasks as ours. Combining appropriate input format and advanced PE, they achieve $\geq$98EM accuracy for 100-digit additions with a 6-layer 8-head model trained on 1–40. Moreover, they achieve a generalizable length of 45 for a model trained on 1–30, 25 for 1–20, and 10 for 1–10 (no length generalization). One big difference between their analysis and ours is they report the maximum accuracy for each testing length over trials, while we report the medians. Thus, we choose a bit lower threshold (95%) for generalizability than theirs. For a better comparison with Zhou et al. (2024b), we report the maximum exact-match (EM) accuracies. See Figures 14 and 15.
Figure 14: Ablation on the trained lengths (1-layer 4-head model trained with position coupling). Here, we report maximum EM accuracies over 8 runs for each tested length.
Figure 15: Ablation on the number of layers (trained with position coupling). Here, we report maximum EM accuracies over 8 runs for each tested length.
B.4 Decimal Integer Addition Task: Comparison & Combination with RoPE
We examine further the possibility of combining our position coupling and RoPE (Su et al., 2024).
RoPE incorporates the positional information into the key and the query vectors by rotating them. Suppose a position ID $m$ is assigned to a key vector ${\bm{k}}$, for every two consecutive entries of ${\bm{k}}$ (i.e., $(k_{2i-1},k_{2i})$), we rotate by a predefined angle $\theta_{i}$ multiplied by $m$:
$\begin{pmatrix}k_{2i-1}\\ k_{2i}\end{pmatrix}\mapsto\begin{pmatrix}k_{2i-1}\cos m\theta_{i}-k_{2i}\sin m\theta_{i}\\ k_{2i-1}\sin m\theta_{i}+k_{2i}\cos m\theta_{i}\end{pmatrix}.$
We also apply a similar rotation to the query vectors (say, ${\bm{q}}$ with position ID $n$). As a result, the attention score for this key-query pair becomes a function of ${\bm{k}}$, ${\bm{q}}$, and the relative distance $n-m$.
Unlike the original implementation of RoPE, we apply rotations based on the re-assigned position IDs according to our position coupling method. We incorporate this RoPE variant into every attention head of every layer. One more difference in implementation is that, during training, we randomly sampled an integer scaler $\ell\in\{1,2,3,4,5\}$ and multiplied it by the rotation angle. By such random re-scaling of rotation angles, we expect the model could handle unseen large rotation angles at test time.
The result on 12-layer models is showcased in Figure 16 (the orange line). Unlike vanilla RoPE (the blue line), which fails immediately outside the trained lengths (1–40), our combination of RoPE and position coupling achieves a much better generalization up to operand lengths 100.
Figure 16: RoPE-based position coupling, 12-layer model.
B.5 Position Coupling for $N\times 2$ Multiplication Tasks
We present an example of the position coupling method for $N\times 2$ ($N$-digit by 2-digit) multiplication, which is omitted from the main text. See Section 6.1 for the experimental results.
B.6 Addition Task with Multiple Summands
The position coupling scheme for the vanilla addition task (with two operands) can naturally extend to the addition task with multiple summands: assign position IDs in ascending order from most significant digits to least significant digits for every operand and the response. Here, we focus on the addition of three summands.
Figure 18: Exact-match accuracy (median over 4 runs) for triple addition task, trained on sequences of length 1-40 with position coupling, NoPE, and random-start APE. For further experiment details, refer to Table 6.
Experiments.
B.7 Position Coupling for Copy/Reverse Tasks
Data Format & Position Coupling.
Each token of the query sequence is a digit (10 distinct characters). We couple the positions in the query and the response by their correspondence. Note that the position ID assigned to the equal token is different for the two tasks because as per our design principle (Sec 3.1), the equal token is grouped to the response tokens and position IDs have to be consecutive numbers within each group.
Figure 19: Illustration of position coupling for copy/reverse tasks.
Experiments.
We compare position coupling with NoPE and random-start APE. We train a model on lengths 1–40 and evaluate its performance on lengths from 5 to 300, at intervals of 5. While a 1-layer 4-head model is used for the position coupling, we observe that the same architecture fails to memorize training samples for both NoPE and random-start APE. Therefore, we use a 6-layer 8-head model for the latter cases as it is commonly used in the literature (Zhou et al., 2024a).
Figure 20: Exact-match accuracy (median over 4 runs) for (a) copying task and (b) reversing task, trained on sequences of length 1–40 with position coupling, NoPE, and random-start APE. For further experiment details, refer to the Table 7.
The experimental results are described in Figure 20. For both copy and reverse tasks, position coupling exhibits near-perfect accuracy across the entire test length (7.5$\times$ for the trained length). In contrast, NoPE and random-start APE immediately fail to length-generalize.
B.8 Position Coupling for Minesweeper Generator Tasks